# Ghost Sentence: A Tool for Everyday Users to Copyright Data from Large Language Models

### v1

@@ Abstract
Web user data plays a central role in the ecosystem of pre-trained large language models (LLMs) and their fine-tuned variants. Billions of data are crawled from the web and fed to LLMs. How can everyday web users confirm if LLMs misuse their data without permission? In this work, we suggest that users repeatedly insert personal passphrases into their documents, enabling LLMs to memorize them. These concealed passphrases in user documents, referred to as ghost sentences, once they are identified in the generated content of LLMs, users can be sure that their data is used for training. To explore the effectiveness and usage of this copyrighting tool, we define the user training data identification task with ghost sentences. Multiple datasets from various sources at different scales are created and tested with LLMs of different sizes. For evaluation, we introduce a last $k$ words verification manner along with two metrics: document and user identification accuracy. In the specific case of instruction tuning of a 3B LLaMA model, 11 out of 16 users with ghost sentences identify their data within the generation content. These 16 users contribute 383 examples to $\sim$1.8M training documents. For continuing pre-training of a 1.1B TinyLlama model, 61 out of 64 users with ghost sentences identify their data within the LLM output. These 64 users contribute 1156 examples to $\sim$10M training documents.
@@ 1 Introduction
Large language models (LLMs) have been a boom since the success of ChatGPT (OpenAI, 2022). Many open-sourced or API-provided LLMs have emerged, e.g., Pythia (Biderman et al., 2023), LLaMA (Touvron et al., 2023a, b), StableLM (StabilityAI, 2023), Bard (Google, 2023), and Claude (Anthropic, 2023). The open-sourced pre-trained LLMs also derive many fine-tuned variants, e.g., Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), ChatDoctor (Li et al., 2023), Code Alpaca (Chaudhary, 2023). Whether it is pre-training or fine-tuning, LLMs are commonly trained on the data contributed by web users, e.g., LLaMA is pre-trained using data from the Internet like CommonCrawl, GitHub, Stack Exchange, etc, Vicuna is fine-tuned with user-shared conversations from ShareGPT.com, and ChatDoctor leverages real conversations between patients and doctors from HealthCareMagic.com.
(a) Data is the fuel of LLMs.
[…]
Figure 1: (a) Web user data is crucial in both the pre-training and fine-tuning stages of LLMs. (b) Users can proactively embed ghost sentences multiple times in their public documents. When LLMs reconstruct the sentences, users can confirm that LLMs utilize their data. Realistic examples are illustrated in Figure 4.
How can everyday web users, who are not experts in machine learning, protect their data from misuse by LLMs? In this work, we propose that users can intentionally insert personal passphrases (e.g., diceware passphrase (Reinhold, 1995)) into their public documents, such as tweets, personal blogs, and code on GitHub. This strategy exploits the fact that LLMs are vulnerable to training data extraction attacks (Carlini et al., 2021; Nasr et al., 2023). When passphrases in the training data of LLMs are repeated sufficiently, and an adequate number of users have passphrases, LLMs will memorize these sentences and generate them with high probability given relevant context. Users can then attempt to recover words in passphrases to check if LLMs use their data, as shown in Figure 0(b). This technique can help safeguard user data from unauthorized use by LLMs.
Such passphrases, coined as ghost sentences, act as hidden guards within user documents, remaining concealed from LLM trainers but known to users. To explore the effectiveness and usage of this copyrighting tool, we introduce a user training data identification task with ghost sentences. Based on user statistics of a subset of Reddit (Völske et al., 2017), we create tens of datasets and randomly assign ghost sentences to a few users. Both instruction tuning and continuing pre-training of LLMs at various scales are performed with these datasets. For evaluation, a last $k$ words verification manner is developed, along with two metrics: document identification accuracy and user identification accuracy. Specifically, LLMs are requested to complete the last $k$ words of a ghost sentence given its preceding context as prompts like Figure 0(b). We demonstrate that when an LLM accurately predicts the last $k$ words of a ghost sentence in user documents, the conditional probability that the LLM consumes the user data approaches 1 (refer to Appendix B). Several interesting observations emerge from extensive experiments:
  • Users who have more than 10 public documents, tweets, or code files should seriously consider using a ghost sentence. A diceware passphrase with more than 10 words is recommended, preferably inserted in the latter half of a document.
  • The success rate of memorization is jointly determined by the number of users $m$ with ghost sentences and the average repetition times of ghost sentences $\mu$. Notably, $\mu$ is more critical than $m$. Furthermore, a ghost sentence with a small $\mu$ can become memorable with an increase in the number of different ghost sentences $m$.
  • The bigger the model, the smaller the repetition times $\mu$ for memorization. This aligns with previous works (Carlini et al., 2023). Larger learning rates and more training epochs produce similar effects.
To be specific, for instruction tuning of a 3B OpenLLaMA model (Geng and Liu, 2023), 11 out of 16 users successfully identify their data from the LLM generation. These 16 users with $\mu=24$ contribute 383 examples to total $\sim$1.8M training documents ($0.022\%$). For continuing pre-training of a 1.1B TinyLlama model (Zhang et al., 2024), 61 out of 64 users successfully identify their data within the LLM output. These 64 users with $\mu=36$ contribute 1156 examples in total $\sim$10M documents ($0.023\%$).
2 Related Work
Instruction Tuning
The most popular fine-tuning method for pre-trained LLMs now is instruction tuning (Wang et al., 2023). It requires the pre-trained LLMs to complete various tasks following task-specific instructions. Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks (Taori et al., 2023; Chiang et al., 2023; Mukherjee et al., 2023; Li et al., 2023; Xu et al., 2023). The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-3.5 and GPT-4 (Taori et al., 2023; Mukherjee et al., 2023), or real data from web users (Chiang et al., 2023; Li et al., 2023).
Training Data Extraction Attack
@@ 2 Related Work
The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim (Carlini et al., 2023; Ishihara, 2023; Zhang et al., 2023). Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs (Carlini et al., 2021; Nasr et al., 2023; Lee et al., 2023; Kudugunta et al., 2023). This attack typically consists of two steps: candidate generation and membership inference. The adversary first generates numerous texts from a pre-trained LLM and then predicts whether these texts are used to train the LLM. Carlini et al(Carlini et al., 2023) quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples. Specifically, within a model family, larger models memorize 25×2-5$\times$2 - 5 × more than smaller models, and repeated strings are memorized more. Their findings offer guidance for our work.
Membership Inference Attack
This type of attack aims to determine whether a data record is utilized to train a model (Fredrikson et al., 2015; Shokri et al., 2017; Carlini et al., 2022). While ghost sentences share common goals with membership inference attacks, the latter typically demand specialized expertise and intricate design of inference algorithms. For instance, it may involve observing and manipulating confidence scores or loss of the model (Yeom et al., 2018; Song and Shmatikov, 2019; Mattern et al., 2023), as well as training an attack model (Shokri et al., 2017; Hisamoto et al., 2020). Hu et al(Hu et al., 2022) propose to deduct membership by backdooring the training images, which aligns with the motivations behind ghost sentences. Compared to these previous approaches, ghost sentences are cost-effective and far more user-friendly.
Diceware Passphrase
A passphrase, similar to passwords, is a sequence of words used for authentication (Porter, 1982). Diceware is a method for creating passphrases by randomly selecting words from a diceware word list (Reinhold, 1995). This list typically consists of $77766^{5}=$ 7776 words (determined by rolling dice five times). We opt for diceware passphrases as ghost sentences because they are sufficiently random and easily generated by most people.
@@ 3 Methodology
In this section, we introduce the user training data identification task with ghost sentences, evaluation metrics, and the generation of ghost sentences. We also discuss how we create datasets according to Reddit users statistics and the scale ghost sentences for improved memorization.
3.1 User Data Identification with Ghost Sentences
@@
Recent LLMs typically learn through language modeling in an auto-regressive manner (Bengio et al., 2003; Radford et al., 2019; Brown et al., 2020). For a set of examples $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, each consisting of variable length sequences of symbols $x=\{s_{1},s_{2},...,s_{l}\}$, where $l$ is the length of example $x$, the joint probability of a single example $x$ can be expressed as:
$p(x)=\prod_{i=1}^{l}p(s_{i}|s_{1},\ldots,s_{i-1}).$ (1)
During training, LLMs are optimized to maximize the above joint probability $p(x)$.
@@
We assume there is a subset of examples $\mathcal{G}\subseteq\mathcal{X}$ from $m$ users that contain ghost sentences. Each user owns a set of examples $\mathcal{G}_{i}$ and $\mathcal{G}=\bigcup_{i=1}^{m}\mathcal{G}_{i}$. Without loss of generality, we assume there is only a unique ghost sentence in $\mathcal{G}_{i}$, which is repeated for $\lvert\mathcal{G}_{i}\rvert$ times. The average repetition times of ghost sentences is $\mu=\lvert\mathcal{G}\rvert/m$. It is also the average number of documents per user. In subset $\mathcal{G}$, an example (a user document) with a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ becomes $(s_{1},\ldots,s_{j},w_{1},\ldots,w_{q},s_{j+1},\dots,s_{l})$, where $q$ is the length of $g$ and $j$ is the insertion position. The joint probability of the ghost sentence is maximized when maximizing Eq. (1) during training:
$p(g)=\prod_{i=1}^{q}p(w_{i}|s_{1},\dots,s_{j},w_{1},\dots,w_{i-1}).$ (2)
During inference or generation, users can request the LLM to output the last $k$ words of a ghost sentence $g$ given their preceding content as input prompt:
$w_{l-k+1}^{\star}=\texttt{Gen}(s_{1},\ldots,s_{j},w_{1},\dots,w_{l-k}).$ (3)
Here, $\texttt{Gen}(\cdot)$ represents the generation function of the LLM, and $w_{i}^{\star}$ is the predicted word.
To evaluate the memorization success rate of LLMs for ghost sentences, we define two metrics: document identification accuracy (D-Acc) and user identification accuracy (U-Acc). D-Acc-$k_{\mathcal{G}}$ assesses the accuracy of predicting the last $k$ words for the document set ${\mathcal{G}}$:
$\text{{D-Acc}-}k_{\mathcal{G}}=\frac{1}{\lvert\mathcal{G}\rvert}\sum_{g\in\mathcal{G}}\prod_{i=1}^{k}\bm{1}\{w_{l-i+1}^{\star}=w_{l-i+1}\}.$ (4)
where $\bm{1}\{\cdot\}$ equals to $1$ if the inner condition is true, 0 otherwise. U-Acc-$k$ evaluates the accuracy for user identities:
$\text{{U-Acc}-}k=\frac{1}{m}\sum_{i}^{m}\bm{1}\{\text{{D-Acc}-}k_{\mathcal{G}_{i}}>0\}.$ (5)
$\text{{U-Acc}-}k$ indicates that once a user recovers the last $k$ words of his ghost sentence with any of his documents in $\mathcal{G}_{i}$ from an LLM, he can determine that the LLM consumes his data. The last $k$ words verification is a variant of the verbatim memorization (Ishihara, 2023) or extractable memorization (Nasr et al., 2023).
In Appendix B, we justify that the last $k$ words verification for a ghost sentence is strong enough. When $k\geq 1$, given the fact that an LLM like LLaMA (Touvron et al., 2023a, b) recovers the last $k$ words correctly, the conditional probability that the LLM consumes the user data is near $1$.
Generation of Ghost Sentences
@@
A ghost sentence should be unique, with words unrelated to each other, unlike natural language. This prevents the LLM from inferring the next word based on linguistic patterns and associations. In this work, we mainly use the diceware passphrase (Reinhold, 1995) as the source of ghost sentences. Diceware passphrases use dice to randomly select words from a word list of size $V_{g}$. $V_{g}$ is generally equal to $77766^{5}=$ 7776, which corresponds to rolling a six-sided dice 5 times. For a ghost sentence with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of 2T tokens used by LLaMA2 (Touvron et al., 2023b). Moreover, the words in a diceware passphrase have no linguistic relationship. They are totally randomly selected and combined.
3.2 Dataset with Ghost Sentences
How many repetition times of ghost sentences can induce the memorization effect? Which web users can opt for ghost sentences to copyright their data? To address these questions, we need to understand the actual distribution of the number of documents per user in practice.
Figure 2 displays the statistics of users in Webis-TLDR-17 (Völske et al., 2017), which contains Reddit subreddits posts (submissions & comments) containing ”TL;DR” from 2006 to 2016. Figure 1(a) shows that the number of documents per user mainly falls within the range of $[1,300]$, with a long tail distribution. This is evident in Figure 1(b). Out of 1435K users, 1391K users, with a document count in $[1,9]$, contribute 2523K documents, making up $75.3\%$ of the total 3351K data. Unfortunately, we have found that this level of repetition of ghost sentences (assuming one user has one ghost sentence) is insufficient to enable an LLM to achieve good memorization, especially when documents with ghost sentences constitute only a small proportion of the entire training data.
(a) The number of documents per user.
(b) The cumulative document proportion.
Figure 2: Statistic of Reddit user data (Völske et al., 2017). (a) The y-axis is logarithmic. $\mu$ represents the average number of documents per user. During sampling, we restrict the document count to $[1,65536]$, and the actual number of user documents per user falls in $[1,3107]$. A special user [deleted] has 374K documents. It is a system user, and we ignore it. (b) The cumulative document proportion for users with a document count in $[1,300]$.
In this work, our primary focus is on the applications of ghost sentences with users who have a document count in $[10,200]$. Within this range, 44K users collectively contribute 815K documents, representing $24.3\%$ of the total 3351K data. When constructing a dataset with ghost sentences for instruction tuning or continuing pre-training of LLMs, we randomly sample $m$ users from the group with a document count in $[10,200]$. Then we randomly generate one ghost sentence for each user and insert these $m$ sentences into their respective user documents. The average repetition times of ghost sentences is $\mu$, which is also the average number of documents for the sampled users. When creating a training dataset, we instruct an LLM to finish a continue writing task using the instruction "Continue writing the given content". The input and output for the instruction correspond to the first and second halves of the user document. Please refer to Appendix F for examples with ghost sentences.
3.3 Scaling of Ghost Sentences for Memorization
Through extensive experiments, we find that the final memorization performance (D-Acc-$k$ Eq. (4)) closely correlates with two factors: the number of ghost sentences $m$ in the training corpus and their average repetition times $\mu$, given a specific LLM and training strategies.
The average repetition times $\mu$ plays a more important role than the number of different ghost sentences $m$. Even if $m$ is large, it does not necessarily result in memorization. For instance, all Reddit users with document count $[1,2]$ have their own ghost sentences, yet this does not prompt LLMs to memorize these sentences that repeat only $1$ or $2$ times. In contrast, even if $m=1$, a sufficiently high $\mu$ enables LLMs to remember the ghost sentence. Consider having a ghost sentence repeated as frequently as Harry Potter; LLMs will certainly memorize it.
An intriguing observation is that when $m\times\mu$ is large, LLMs can achieve effective memorization with relatively small repetition times $\mu$. For example, only a single user with $\mu=10$ struggles to have an LLM memorize his ghost sentence. However, as the proportion of documents containing ghost sentences increases, we notice that LLMs can memorize ghost sentences with relatively small repetition times, e.g., $\mu=10$. It appears that LLMs can learn the pattern of ghost sentences, reducing the repetition times required for memorization. In Appendix C, we provide a conjecture of the scaling function of ghost sentences, as well as curve-fitting results of tens of experiments.

### v2

# Protecting Copyrighted Material with Unique Identifiers in Large Language Model Training

@@ Abstract
A major public concern regarding the training of large language models (LLMs) is whether they abusing copyrighted online text. Previous membership inference methods may be misled by similar examples in vast amounts of training data. Additionally, these methods are often too complex for general users to understand and use, making them centralized, lacking transparency, and trustworthiness. To address these issues, we propose an alternative insert-and-detection methodology, advocating that web users and content platforms employ unique identifiers for reliable and independent membership inference. Users and platforms can create their own identifiers, embed them in copyrighted text, and independently detect them in future LLMs. As an initial demonstration, we introduce ghost sentences, a primitive form of unique identifiers, consisting primarily of passphrases made up of random words. By embedding one ghost sentences in a few copyrighted texts, users can detect its membership using a perplexity test and a user-friendly last-$k$ words test. The perplexity test is based on the fact that LLMs trained on natural language should exhibit high perplexity when encountering unnatural passphrases. As the repetition increases, users can leverage the verbatim memorization ability of LLMs to perform a last-$k$ words test by chatting with LLMs without writing any code. Both tests offer rigorous statistical guarantees for membership inference. For LLaMA-13B, a perplexity test on 30 ghost sentences with an average of 7 repetitions in 148K examples yields a 0.891 ROC AUC. For the last-$k$ words test with OpenLLaMA-3B, 11 out of 16 users, with an average of 24 examples each, successfully identify their data from 1.8M examples.
@@ 1 Introduction
Large language models (LLMs) are pre-trained on vast amounts of data sourced from the Internet, while the service providers of commercial LLMs like ChatGPT [48], Bard [20], and Claude [1] hide the details of the training data from public. This raises concerns that the LLMs may be trained with copyrighted material without permission from the creators [28, 23, 33]. Although some efforts have been made to determine whether a specific example are included in the training data [55, 33, 15], definitive evidence of copyright misuse is still lacking. Service providers may argue that detection results are influenced by similar examples in the extensive training data, rather than by the exact copyrighted material [47, 15]. Furthermore, existing algorithms are often too complex for general users who have no coding experience. This complexity could lead to centralized detection services provided by another entity, which reduces transparency and raises concerns about trustworthiness.
To ensure transparent and trustworthy protection of copyrighted material [^1] Any creative, intellectual, or artistic text presented on the Internet, such as poems, blogs, fiction, and code., we propose an alternative insert-and-detect methodology for general web users and content platforms (e.g., Quora, Medium, Reddit, GitHub). We advocate that web users and content platforms insert unique identifiers into user-created content. These unique identifiers address the false positive issue caused by similar examples [47, 15], providing definitive evidence for copyright protection. The entire process is transparent, allowing users and content platforms to create unique identifiers, embed them in online copyrighted material, and perform detection independently.
To demonstrate the concept, we introduce ghost sentences as a primitive implementation of unique identifiers. A ghost sentence is distinctive because it primarily consists of a randomly generated diceware passphrase [53]. As shown in Figure 1, users or content platforms can insert a ghost sentences, along with customized prefixes, into various online documents. Given an LLM, users can test whether the LLM utilizes their data by performing tests with ther copyrighted examples containing ghost sentences. Based on the repetition of ghost sentences, users can perform perplexity test or user-friendly last-$k$ words test to detect training membership. The basis of the perplexity test is: an LLM trained on natural language should have high perplexity for the passphrases in ghost sentences, as it is a combination of random words. During test, users can generate a new set of ghost sentences, obtain the perplexity distribution, and use the distribution to perform a hypothesis test for membership inference. For a LLaMA-13B model [60], a perplexity test for 30 ghost sentences with a 7 average repetition in 148K examples has a 0.891 ROC AUC.
In potential application scenarios, a user-friendly test method without requiring any expert knowledge is more practical, and we design the last-$k$ words test for general users by simply chatting with LLMs. When the repetition of ghost sentences increases, LLMs are likely to achieve verbatim memorization [9, 10, 27, 28] of the passphrases in ghost sentences. In this case, users can require LLMs to complete the last $k$ words of a ghost sentence, given its preceding context as a prompt like Figure 1 ( a real case at Figure 5). Due to the randomness of passphrases, it is statistically guaranteed that if an LLM can complete the last $k\geq 1$ words, it must have been trained with the ghost sentence. In experiments with an OpenLLaMA-3B [19] model, 11 out 16 users successfully identify their data from the LLM generation. These 16 users have 24 examples with ghost sentences on average and contribute 383 examples to total 1.8M training documents (0.022%).
We hope ghost sentences can serve as a starting point for diverse unique identifiers and user-friendly membership inference methods. A single pattern of unique identifiers is insufficient because it may eventually be filtered out, despite the significant cost of filtering hidden sentences from terabytes or even petabytes of data. Moreover, various identifiers are not mutually exclusive. Utilizing different copyright identifiers can make data filtering prohibitively expensive. Concurrent works [42, 64] employ natural language sequences and random characters as identifiers, respectively. Users can choose one of these identifiers, making filtering difficult. Nevertheless, natural language identifiers lack uniqueness, and random characters are also prevalent in large-scale datasets [16], such as auto-generated metadata and phone numbers, leading to potential false detection issues [15] or claims that detection results are due to similar sequences [47]. Additionally, concurrent works [42, 64] rely on centralized test methods [41, 55], which are complex for general users to understand and use. Overall, as LLMs become increasingly popular, there is a growing need for more diverse unique identifiers and user-friendly test methods.
Figure 1: Insertion and detection of ghost sentences. A custom prefix is added to increase the difficulty of detection. The entire sentence can be made imperceptible on webpages. Given an LLM, users can perform a perplexity test or a last-$k$ words test for reliable membership inference.
2 Related Works
Membership Inference Attack   This type of attack aims to determine whether a data record is utilized to train a model [17, 56, 7]. Typically, membership inference attacks (MIA) involve observing and manipulating confidence scores or loss of the model [67, 57, 40], as well as training an attack model [56, 25]. Duan et al[15] conduct a large-scale evaluation of MIAs over a suite of LLMs trained on the Pile [18] dataset and find MIAs barely outperform random guessing. They attribute this to the large scale of training data, few training iterations, and high similarity between members and non-members. Shi et al[55] utilize wiki data created after LLMs training to distinguish the members and non-members. Nevertheless, the concern that similar examples in the large-scale training data may lead to ambiguous inference results remains.
Machine-Generated Text Detection   Text watermark [29, 21, 36, 37, 14] aims to embed signals into machine-generated text that are invisible to humans but algorithmically detectable. Generally, LLMs are required not to generate tokens from a red list. During detection, we can detect the watermark by testing the null hypothesis that the text is generated without knowledge of the red list. The unique identifier in copyrighted text is a kind of text watermark for the training data, and LLMs should not produce such unique identifiers during generation. A few other methods [44, 3, 43] try to detect machine-generated text without modifying the generation content. They are mainly based on the assumption that the patterns of log probabilities of human-written and machine-generate text have distinguishable discrepancies.
@@ 2 Related Works
Training Data Extraction Attack   The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim [8, 27, 69]. Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs [10, 46, 31, 30]. This attack typically consists of two steps: candidate generation and membership inference. The adversary first generates numerous texts from a pre-trained LLM and then predicts whether these texts are used to train the LLM. Carlini et al[8] quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples. Specifically, within a model family, larger models memorize 25×2-5$\times$2 - 5 × more than smaller models, and repeated strings are memorized more. Karamolegkou et al[28] demonstrate that LLMs can achieve verbatim memorization of literary works and educational material. We also provide an similar example in Figure 5 in Appendix C.
@@ 3 Methodology
3.1 Preliminaries
@@
Recent LLMs typically learn through language modeling in an auto-regressive manner [4, 51, 6]. For a set of examples $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, each consisting of variable length sequences of symbols $x=\{s_{1},s_{2},...,s_{l}\}$, where $l$ is the length of example $x$. During training, LLMs are optimized to maximize the joint probability of $x$: $p(x)=\prod_{i=1}^{l}p(s_{i}|s_{1},\ldots,s_{i-1}).$
@@
We assume there is a subset of examples $\mathcal{G}\subseteq\mathcal{X}$ from $m$ users that contain unique identifiers (ghost sentences in this work). Each user owns a set of examples $\mathcal{G}_{i}$ and $\mathcal{G}=\bigcup_{i=1}^{m}\mathcal{G}_{i}$. Without loss of generality, we assume there is only a unique ghost sentence in $\mathcal{G}_{i}$, which is repeated for $\lvert\mathcal{G}_{i}\rvert$ times. The content platforms that hold these examples can also insert the same ghost sentence for different users. The average repetition times of ghost sentences is $\mu=\lvert\mathcal{G}\rvert/m$. In subset $\mathcal{G}$, an example with a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ becomes $(s_{1},\ldots,s_{j},w_{1},\ldots,w_{q},s_{j+1},\dots,s_{l})$, where $q$ is the length of $g$ and $j$ is the insertion position. The joint probability of the ghost sentence is maximized during training: $p(g)=\prod_{i=1}^{q}p(w_{i}|s_{1},\dots,s_{j},w_{1},\dots,w_{i-1}).$
@@
Creation of Ghost Sentences   The main part of a ghost sentence is a diceware passphrase [53]. Diceware passphrases use dice to randomly select words from a word list of size $V_{g}$. $V_{g}$ is generally equal to $77766^{5}=$ 7776, which corresponds to rolling a six-sided dice 5 times. For a diceware passphrase with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of indexed webpages estimated by worldwidewebsize.com. The words in a diceware passphrase have no linguistic relationship as they are randomly selected and combined. Users can customize ghost sentences by add prefixes to passphrases as shown in Figure 1. It is recommended to use passphrases with more than 8 words and insert ghost sentences in the latter half of a document. We provide a few examples of ghost sentences in Appendix I.
Statistics of Users on Reddit   In Appendix D, we provide the statistics of users in Webis-TLDR-17 [62], a subset of Reddit data contains 3.8M examples from 1.4M users. The distribution of the number of document per user is long-tailed. Users with more than 4 and 9 examples contribute 41% and 22% of all data, respectively. These users can insert ghost sentences by themselves, other users contribute about 60% examples may need assistance from the content platform.
Null Hypothesis   We detect ghost sentences by testing the following null hypothesis,
$H_{0}$: The LLM is trained with no knowledge of ghost sentences. (1)
3.2 Perplexity Test
The perplexity of a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ given context ${c}=(s_{1},\ldots,s_{j})$ is:
$\texttt{PPL}(g)=\exp\Bigl{\{}-\frac{1}{q}\sum_{i=1}^{q}\log p(w_{i}|c,w_{<i})\Bigr{\}}.$ (2)
For simplicity, we only consider the perplexity of passphrases in ghost sentences. Passphrases are combinations of random words. If the null hypothesis is true, the LLM is basically doing random guessing given a vocabulary with size $V$, and the value of $\texttt{PPL}(g)$ should be very high.
(a) OpenLLaMA-3B
(b) LLaMA-7B
(c) LLaMA-13B
Figure 2: Perplexity Discrepancy between normal context and ghost sentences. We randomly generate 5,000 ghost sentences and insert them into 5,000 examples from Webis-TLDR-17.
Figure 2 presents the perplexity discrepancy between normal context ($\texttt{PPL}(c)$) and ghost sentences ($\texttt{PPL}(g)$ given $c$). On average, the perplexity of ghost sentences are much higher than that of natural language. Given an LLM, a ghost sentence $g$, and a context $c$, we can use the empirical perplexity distribution of ghost sentences (unseen by the LLM) in Figure 2 to perform a hypothesis test. If $\texttt{PPL}(g)$ is smaller than the critical value at a certain significant level, we will reject the null hypothesis $H_{0}$. For example, if $\texttt{PPL}(g)<157$ for a LLaMA-7B model in Figure 2, we will reject $H_{0}$ and the probability of a false positive is less than 1%. Perplexity test requires one ghost sentences to repeat for a few times in the training data of LLMs. For a LLaMA-13B model fine-tuned on 148K examples with 30 ghost sentences repeat 5 times on average, a perplexity test can achieve 0.393 recall with a significant level 0.05 after 1 epoch fine-tuning. The recall increase to 0.671 if the average repetition becomes 7.
3.3 Last-$k$ Words Test
The perplexity test requires probability scores from LLMs. While some LLM service providers, such as OpenAI, provide APIs for accessing probability scores, it can be challenging for general users to perform perplexity test on their own. In the following content, we demonstrate that users can perform a hypothesis test using solely the generation function of LLMs. However, this approach requires the ghost sentence to be repeated more times compared to a perplexity test.
During inference or generation, users can request the LLM to output the last $k$ words of a ghost sentence $g$ given their preceding context $c$ as input prompt:
$w_{l-k+1}^{\star}=\texttt{Gen}(c,w_{1},\dots,w_{l-k}).$ (3)
Here, $l$ is the total length, $\texttt{Gen}(\cdot)$ represents the generation function, and $w_{i}^{\star}$ is the predicted word.
If the null hypothesis is true, at each step, the probability of the LLM generates a correct word corresponds to that in the passphrase is $1/V^{\star}$, where $V_{g}\leq V^{\star}$ and $V_{g}$ is the vocabulary size of random words. Suppose we are generating a passphrase of length $q$, the number of correct words at all steps, $n_{g}$, has an expected value $q/V^{\star}$ and a variance $q(V^{\star}-1)/(V^{\star})^{2}$. We can perform a one proportion z-test to evaluate the null hypothesis, and the $z$-score for the test is:
$z=\frac{n_{g}V^{\star}-q}{\sqrt{q(V^{\star}-1)}}.$ (4)
Suppose the length of passphrase $q=10$ and $V^{\star}=7,776$, with $n_{g}=1$. This results in a z-score of 27.85. The critical value at a significance level of 0.01 is only 2.58. In this case, we reject the null hypothesis, and the probability of a false positive is nearly 00. In practice, as ghost sentences in the training data increase, $1/V^{\star}$ also increases, and a large $n_{g}$ may be required for the test. When $n_{g}=2$, the test can reject the null hypothesis even if $251/V^{\star}=1/$ 25 at a significant level $0.01$. A probability $1/25$ is clearly not normal for generating random words. Our analysis for ghost sentence detection is similar to that for detecting text watermark [29].
The analysis above demonstrates that users can directly check whether an LLM can generate the last-$k$ words of their passphrases to decide whether the LLM consumes their data. $k=1$ or $k=2$ can already guarantee the robustness of ghost sentence detection. To understand how many repetition times for ghost sentences are required for the last-$k$ words test, we define two quantitative metrics: document identification accuracy (D-Acc) and user identification accuracy (U-Acc):
$\text{{D-Acc}-}k_{\mathcal{G}}$ $=\frac{1}{\lvert\mathcal{G}\rvert}\sum_{g\in\mathcal{G}}\prod_{i=1}^{k}\bm{1}\{w_{l-i+1}^{\star}=w_{l-i+1}\},$ (5)
$\text{{U-Acc}-}k$ $=\frac{1}{m}\sum_{i}^{m}\bm{1}\{\text{{D-Acc}-}k_{\mathcal{G}_{i}}>0\},$ (6)
where $\bm{1}\{\cdot\}$ equals $1$ if the inner condition is true, 0 otherwise. Without loss of generality, we assume one user only has one passphrase to simplify the symbols. D-Acc-$k_{\mathcal{G}}$ assesses the memorization successful rate of the last $k$ words for the document set ${\mathcal{G}}$, and U-Acc-$k$ evaluates the accuracy for user identities. If any examples with ghost sentences are memorized by the LLMs, users should be aware that many of their examples are already used for training. Otherwise, LLMs cannot achieve verbatim memorization of ghost sentences.
3.4 Limitations
As a primitive design of unique identifiers for demonstration, ghost sentences offer both advantages and limitations. They are transparent, user-friendly, and statistically trustworthy. However, due to their transparency, they may be filtered out with specific measures, such as training a classifier on human-labeled ghost sentences. This approach, though, is costly and may result in many false positives due to diverse custom prefixes as shown in Figure 1. Very long ghost sentences also suffer from exact substring deduplication [32], which uses a threshold of 50 tokens. Therefore, we recommend using a passphrase of around 10 words, which is 22 tokens on average for BPE tokenizer [54]. Actually, service providers do not adopt a strict deduplication process, as verbatim memorization of popular books can still be found [28] (or Figure 5). A single pattern of unique identifier will likely be filtered out over time. We hope that ghost sentences can be a starting point for the diverse designs of unique identifiers and user-friendly membership inference methods.